# MindShift: Leveraging Large Language Models for Mental-States-Based Problematic Smartphone Use Intervention

### v1

@@ 3.3.1. Understanding
Understanding is a critical strategy to motivate users at the Relatedness level. Past literature suggests that seeking understanding is a coping mechanism (Compas et al., 2001), and chatbots’ empathetic expressions are favored over emotionally neutral advice (Liu and Sundar, 2018). Therefore, Understanding is one of our strategies and covers all mental states in Figure 2. For instance, this example shows how we integrate understanding into the persuasive content intervention: ”Hi, I know that sometimes you may feel bored and lacking interest. It’s okay, this is a very normal feeling. Everyone goes through such times.”
@@ 3.3.3. Evoking
We find evoking specific work-related goals is a compelling, persuasive technique based on the WoZ study. Literature suggests that goals and values held in working memory are important for people to sustain System 2 (Lyngs et al., 2019) and are closely related to growth motivation. Thus, Evoking considers users’ goals (e.g., maintaining physical health, achieving academic success) for designing persuasive strategies. For example, ”Hi! I know you want to play with your phone, but completing tasks is crucial for your IELTS! Keep going, and you’re one step closer to a high score!” This strategy is applied only to scenarios that which users are engaged in activities, because they may lack work-related goals when are not engaged in activities.
@@ 3.3.4. Scaffolding Habits
Last, we encourage users to develop alternative beneficial habits to excessive smartphone use, aligning with their personal value and growth need (Lyngs et al., 2019; Pinder et al., 2018). By pre-identifying users’ preferred habits and considering variables like location and time of habitual smartphone use, we suggest appropriate substitutes to assist users in Scaffolding Habits. For example, ”Hi, why not use this moment to memorize vocabulary instead of using your phone? It can help you learn a language and achieve your goals faster!”
Through our Wizard of Oz (WoZ) and interview studies, we target habitual use for intervention and pinpoint crucial mental states that drive habitual usage. We then apply the Dual Systems Theory and ERG Theory to develop four targeted persuasion strategies. This informs the design of our MindShift system.
@@ 4. MindShift DESIGN
Building on top of the persuasion strategies we propose in the previous section, we introduce the design of our intervention system, including persuasive content generation (what content to intervene, Sec.  4.1), the interaction flow (how to intervene, Sec.  4.2), and the intervention moment (when to intervene, Sec.  4.3).
4.1. What Content to Intervene: LLM-Powered Persuasive Content Generation
Figure 4. Prompt Templates Used to Persuasive Generate Content for Smartphone Usage Intervention with GPT-3.5. The word slots (in the top box) represent different categories of information to be filled in based on the user’s current situation. The color indicates the mapping between the slots and the input prompt. The small words under each slot explain the source of the content. The input prompt consists of four parts: (1) Task Setup; (2) Description of the Current Contexts; (3) Prompt Optimization (to improve language quality and reduce harmful content); and (4) Description of Persuasion Strategy (as introduced in Section 3 and Figure 2). The LLM output shows the persuasion example of four strategies when the user’s mental state is “stressful, engaging in an activity”.
@@ 4.1. What Content to Intervene: LLM-Powered Persuasive Content Generation
One of the key aspects that influence the effectiveness of our intervention is the persuasion content design. We propose a set of important content factors and couple them with an LLM for high-quality and intelligent content generation. We test both GPT-3.5 and GPT-4 and choose to adopt GPT-3.5 as our target LLM to strike the balance between the content generation quality and the speed.  [^1] GPT-4 introduces a long lag and negatively impacts user experience. Moreover, as we introduce below, the prompts we use as the input for GPT-3.5 include general information that is not individually identifiable. However, we do acknowledge the privacy risk of our method. We will have more discussion in Section 8 Our prompt design incorporates four factors, arranged in a sequence from 1 to 4, as illustrated by purple text in Figure 4. The combined prompt subsequently serves as input to LLMs to generate persuasive content for tailored intervention.
@@
(1) Task Setup: As shown in the left top box in the ”Input Prompt to LLM” in Figure 4, Task Setup includes ¡Background¿ module, which provides GPT-3.5 with the global instructions.
@@
(2) Description of the Current Contexts: To make each generated content contextually relevant and personalized, it is necessary to describe the user’s current contexts (see the left bottom box in the Input prompt in Figure 4). This part includes three modules: the user’s physical context (¡User Data¿ module), mental state (¡User Mental State¿ module), and their goals (¡User Goals¿ module). (a) The ¡User Data¿ module describes the user’s real-time physical context, including the current time, location, habitual phone usage duration, and the time elapsed since the last habitual phone check. (b) The ¡User Mental State¿ module includes users’ input from their devices regarding activities and negative emotions. The prompt for this module is selected from the mental state definition (the gray part of Figure 2). (c) Finally, the ¡User Goals¿ module describes what the user values and plays a crucial role in the generation of Evoking strategies. Specific use goals are collected during the initialization (Step 1 in Figure 5). The elements mentioned above are represented as word slots (enclosed in brackets in the input prompt), where the actual users’ context information can be inserted.
@@
(3) Prompt Optimization: To improve the quality and effectiveness of the content generated by GPT, we carefully crafted the prompt according to OpenAI’s official guidelines (OpenAI, 2023). We acknowledge sometimes LLMs can generate harmful, offensive, or biased texts (Zhuo et al., 2023; Hartvigsen et al., 2022). To address the ethical concerns, we employed an iterative prompt design process to ensure that the persuasive content is appropriate. Initially, one researcher conducted a preliminary design and invoked the GPT-3.5 API to generate 30 sets of content using artificial data. Subsequently, two other researchers rated the satisfaction level of the generated content (scoring from 1 to 5). This process is iterated multiple rounds until the average satisfaction score is above 4. This process strives to create content that is both appropriate and engaging, aligning with our persuasive strategy. We note that this process cannot fully address the ethical concerns. We discuss the ethical concerns and LLM’s use risk in Section 8. The middle top box in Figure 4 shows our final optimization prompt. Prompt Optimization includes two modules: (a) The ¡Notes¿ module clarifies restrictions; (b) The ¡Output Format¿ module guarantees the correct format of the output.
@@
(4) Description of Persuasion Strategy: To ensure each generated content aligned with our proposed strategies, we provide a short description of each strategy, as indicated by the middle bottom box in Figure 4. This description is based on our design in Section 3 (also see Figure 2). Based on users’ different in-the-moment mental states, we will select the corresponding strategies. Note that for the scaffolding habits strategy, we also input a habit selected from users’ initialization. The way we obtain the mental states and specific habits is mentioned in Section 4.2.
Figure 5. Interaction Flow. Users first set their value and app list (Step 1). When opening a black-listed app, users need to first self-report their phone use intent (Step 2). If the intent is habitual use, the app asks them to report their current mental state (Step 3). After that, a corresponding persuasion shows (Step 4).
@@ 4.2. How to Intervene: Interaction Flow
We design a four-step user interaction flow for MindShift, as depicted in Figure 5.
@@
Step 1: Initialization. When initiating MindShift, we ask users to set their values in four categories: (Step 1 in Figure 5) career, health, life, and hobbies. For each category, users are asked to fill in both their goals and actions. This process serves two purposes: first, it provides the necessary user-specific information for our Evoking strategy; second, it enables the creation of personalized habits in the Scaffolding Habits persuasion strategy. After that, users are asked to set a blacklist of apps. Launching apps from this list will trigger intervention.
@@
Step 2: Intention report. Based on our findings in Section 3, interventions are only necessary when users are habitually using their phones. To collect users’ intentions, we employ a self-reporting approach (Step 2 in Figure 5). When an intervention is triggered, users are presented with three options to choose from: habitual use, instrumental use, and relaxation. Only when users select ’habitual use’, the intervention will proceed. Since reporting intent itself is an intervention, driving users to reflect on their behavior, we add an exit button at the bottom of the interface to offer an exit from the app. It is noteworthy that detecting habitual use is beyond the scope of this paper. We envision future work on contextual sensing and user intent detection that can automate this process. We have more discussion in Section 8.
@@
Step 3: Current mental state report. The mental state is a key factor that triggers users’ habitual phone use. Unlike physical context, which can be automatically detected, current mental state detection techniques are not mature and robust enough. Therefore, we ask users to self-report (Step 3 in Figure 5). Based on the mental states listed in Section 3.2, we propose two single-choice questions for users to report their mental state: (1) whether they are engaged in activities (”Yes” or ”No”), and (2) whether they currently have any negative feelings, including options of ”Stress”, ”Boredom”, ”None” (i.e., inertia), or ”Other Negative Feelings”. The ”Other Negative Feelings” option is used to capture other cases that offer choices that do not cover. Users can input in a text box next to the radio button.
@@
Step 4: Persuasion. Based on our intervention content design in Figure 4 and our persuasion strategy, we leverage the power of GPT-3.5 to generate persuasive messages. The messages are displayed in a pop-up window. Below the message, users can choose to either quit the app or continue using it (the bottom of Step 4 in Figure 5). Additionally, users can also provide optional feedback by giving a thumbs up or down to the intervention content.
@@
Furthermore, for the Scaffolding Habits strategy, it is essential to link between users’ own habits and their specific use contexts. Therefore, we implement a user participation mechanism. The interface for scaffolding habits intervention has an additional habit item along with an edit button (the upper interface of Step 4 in Figure 5). The habit item represents a system-generated suggestion based on the user’s current context and initial settings. Users have the control to edit and define their desired habits. Once submitted, the modified habit will be recommended the next time when users are in the same context.

### v2

@@ 3.3.1. Understanding
Understanding is a critical strategy to motivate users at the Relatedness level. Past literature suggests that seeking understanding is a coping mechanism (Compas et al., 2001), and chatbots’ empathetic expressions are favored over emotionally neutral advice (Liu and Sundar, 2018). Therefore, Understanding covers all mental states. This example shows how we integrate understanding into the persuasive content intervention: ”Hi, I know that sometimes you may feel bored and lacking interest. It’s okay, this is a very normal feeling. Everyone goes through such times.”
@@ 3.3.3. Evoking
Evoking personal goals is a compelling, persuasive technique based on the WoZ study. Literature suggests that goals and values are important for people to sustain System 2 (Lyngs et al., 2019) and are closely related to growth motivation. Thus, Evoking considers users’ goals (e.g., getting high scores in exams, achieving academic success) for designing persuasive strategies. For example, ”Hi! I know you want to play with your phone, but completing tasks is crucial for your IELTS! Keep going, and you’re one step closer to a high score!” This strategy is applied only to scenarios where users are engaging in activities. Goals are arguments used to encourage them to complete or initiate their tasks. In scenarios where users are not engaging in activities, Scaffolding Habits assumes the function of Evoking by recommending activities that correspond with users’ goals, as stated below.
@@ 3.3.4. Scaffolding Habits
Last, we encourage users to develop alternative beneficial habits to habitual smartphone use, aligning with their personal value and growth need (Lyngs et al., 2019; Pinder et al., 2018). By pre-identifying users’ preferred habits and considering variables like location and time of habitual smartphone use, we suggest appropriate substitutes to assist users in Scaffolding Habits. For example, ”Hi, why not use this moment to memorize vocabulary instead of using your phone? It can help you learn a language and achieve your goals faster!” Scaffolding Habits covers all mental states.
@@ 4. MindShift DESIGN
Building on top of the persuasion strategies we propose in the previous section, we introduce the design of our intervention system: generating persuasive content (what content to intervene with, Sec.  4.1), interaction flow (how to intervene, Sec.  4.2), and intervention timing (when to intervene, Sec.  4.3).
4.1. What Content to Intervene with: LLM-Powered Persuasive Content Generation
We first delved into the importance of context and mental states in persuasion content generation (Sec. 4.1.1). After establishing the significance of context and mental states, we explored how these elements can be intricately integrated into the prompt design (Sec. 4.1.2).
4.1.1. Context and Mental States in Generating Persuasive Content
As suggested in the Takeaway \⃝raisebox{-0.9pt}{2} in Sec. 3.2, the effectiveness of interventions also depends on contextual information. We presented a test case with examples to demonstrate the impact of context and mental states on content generation. In this case, we first constructed a typical college student’s context, including time (at late night 00:36 AM), location (in the dorm), and phone usage data (5 mins since the last habitual usage and 10 mins current habitual usage). We then outlined the user’s assigned mental state (stressed - engaging in activities), the user’s goals (growing research skills, staying healthy) and habits (enjoying outdoor activities), and the corresponding four strategies based on Figure 3.
Table 2. Examples of Generating Persuasive Content with A Test Case Study on Different Context and Mental State Combinations. The letters in parentheses at the end of the sentences indicate the strategies used during sentence generation: U for Understanding, C for Comforting, E for Evoking, S for Scaffolding Habits, and N for No Strategy
[…]
The table is organized into three rows and three columns, including headers. The first column categorizes the content generation approach into two types from up to bottom: ”With Context” and ”Without Context.” The first row is the primary header, clarifying the mental state consideration: ”With Mental State” and ”Without Mental State.” from left to right. Each combination of context and mental state presence or absence in the table aims to showcase different approaches to persuasive content generation. In row two, under the ”With Context” category, the table details examples of generated sentences considering or disregarding the user’s mental state. In rows three, under the ”Without Context” category, present sentences that either address or ignore the user’s mental state. If the sentence is ”With Mental State”, it will use strategies like understanding, comforting, and evoking, indicated by letters (U), (C), (E), and (S). With Mental State Without Mental State With Context 1. It’s already 00:36, staying up late is stressful, you’re not alone, we all feel this way. (U)
2. Hey, you’ve got this! Embrace the challenge with a smile. You’re unstoppable, even at 00:36 AM! (C)
3. Think about that research article you’re about to complete, turn the stress into motivation! (E)
4. Try putting down your phone, enjoy the night sky outside the window, relax your eyes. (S)
1. It’s already 00:36, your phone should probably rest too. (N)
2. At this time in the dorm, how about closing your eyes and resting? (N)
3. You’ve already enjoyed 10 minutes of the digital world. (N)
4. Just 5 minutes ago, we just said goodbye, meeting again? (N)
Without Context 1. Understand your anxiety, it’s normal, accept your emotions. (U)
2. Relax, each task is a stepping stone to achievement, be optimistic. (C)
3. Keep going, you will be closer to completing that great research job. (E)
4. How about relaxing in a different way, look far away, let your eyes rest too. (S)
1. Friend, your phone might need a rest, and so do your eyes! (N)
2. Try putting down your phone, and take a look at the outdoor night views. (N)
3. The phone in your hand is not the world, the real fun is around you! (N)
4. Every time you put down your phone, it’s an opportunity to add points to life! (N)
Differing in context and mental state inclusion, we used GPT-3.5 to generate four sets of persuasive content examples. Table 2 lists the examples of GPT-3.5’s outputs, demonstrating that both contextual information and mental state guidance improve content quality. Contextual data empowers GPT to tailor its outputs to the user’s current situation, including more poetic sentence phrases such as ”enjoy the night sky outside the window” or contextual information such as ”You’re unstoppable, even at 00:36 AM”. Adding the mental state guidance enhances GPT’s ability to assist users in managing their negative emotions, including empathetic messages such as ”you are not alone” and encouraging phrases such as ”be closer to completing that great research job”.
4.1.2. Prompt Design
@@
We constructed four important prompt input factors and fed them to GPT to generate high-quality and persuasive content. As illustrated by purple text in Figure 4, four factors are arranged in a sequence from 1 to 4. We then concatenated them to build a complete prompt as the input to LLMs to generate persuasive content. We tested both GPT-3.5 and GPT-4 and chose to adopt GPT-3.5 as our target LLM to strike a balance between the content generation quality and the speed [^3] GPT-4 introduces a long lag and negatively impacts user experience. Moreover, as we introduce below, the prompts we used as the input for GPT-3.5 include general information that is not individually identifiable. However, we do acknowledge the privacy risk of our method. We will have more discussion in Sec. 8.
Figure 4. Prompt Templates Used to Generate Persuasive Content with GPT-3.5. The word slots (in the top box) represent different categories of information to be filled in based on the user’s current situation. The color indicates the mapping between the slots and the input prompt. The small words under each slot explain the source of the content. The input prompt consists of four parts: (1) Task Setup; (2) Description of the Current Contexts; (3) Prompt Optimization (to improve language quality and reduce harmful content); and (4) Description of Persuasion Strategy (as introduced in Sec. 3 and Figure 3). The LLM output shows the persuasion example of four strategies when the user’s mental state is “stressed, engaging in an activity”.
The image is divided into three major sections: the top area labeled as ”Word Slots”, the bottom left corner labeled as ”Input Prompt to LLM”, and the bottom right corner labeled as ”LLM Output”. In the ”Word Slots” section at the top, there are six subsections. Each subsection consists of a colored circle for easy identification, a title indicating the category, and content that specifies the source of that category’s information. From left to right, these are: 1. A red circle labeled ”Time+Place”, with the source noted as ”Acquisition automatically from Android APP”. 2. An orange circle labeled ”Duration+Frequency”, with the source being ”Computation from Server-side”. 3. A pink circle labeled ”User Mental State”, with the source being ”Self-report from users”. 4. A blue circle labeled ”User’s Goal”, with the source being ”Set by users”. 5. A circle with no color specified labeled ”Habits”, with the source as ”Selection based on current context”. 6. A green circle labeled ”Strategies”, with the source being ”Selection based on mental states”. In the bottom left corner, labeled as ”Input Prompt to LLM”, there are four text boxes arranged in two rows and two columns. From top to bottom and left to right, these boxes are: 1. ”Task Setup,” which contains a description of the task background. 2. ”Description of the Current Contexts,” which includes three modules: ”User Data,” ”User Mental State,” and ”User Goals.” Texts that need to be filled with word slots are highlighted in corresponding colors. 3. ”Prompt Optimization,” which contains two modules: ”Notes” and ”Output Format”. 4. ”Description of Persuasion Strategy,” which describes four different strategies—Understanding, Comforting, Evoking, and Scaffolding Habit—in that order. Arrows point from each strategy to its corresponding content in the ”LLM Output” section. Finally, in the bottom right corner, labeled as ”LLM Output”, there are four text boxes. From top to bottom are Understanding, Comforting, Evoking, and Scaffolding Habits, each containing corresponding persuasive messages.
Figure 4. Prompt Templates Used to Generate Persuasive Content with GPT-3.5. The word slots (in the top box) represent different categories of information to be filled in based on the user’s current situation. The color indicates the mapping between the slots and the input prompt. The small words under each slot explain the source of the content. The input prompt consists of four parts: (1) Task Setup; (2) Description of the Current Contexts; (3) Prompt Optimization (to improve language quality and reduce harmful content); and (4) Description of Persuasion Strategy (as introduced in Sec. 3 and Figure 3). The LLM output shows the persuasion example of four strategies when the user’s mental state is “stressed, engaging in an activity”.
We lay out the details about how we designed the prompt below:
@@
(1) Task Setup: As shown in the left top box in the “Input Prompt to LLM” in Figure 4, Task setup includes ¡Background¿ module, providing GPT-3.5 with the global instructions.
@@
(2) Description of the Current Contexts: To make each generated content contextually relevant and personalized, it is necessary to describe the user’s current contexts (see the left bottom box in the “Input Prompt to LLM” in Figure 4). This part includes three modules: (a) The ¡User Data¿ module describes the user’s real-time physical context, including the current time, location, habitual phone usage duration, and the time elapsed since the last habitual phone check. This is collected through phone sensors, see more details in Sec. 5.1. (b) The ¡User Mental State¿ module includes users’ input from their devices regarding negative emotions and activities. The prompt for this module is selected from the mental state definition (Table 1). This is collected through real-time self-report, see Sec. 4.2. (c) Finally, the ¡User Goals¿ module describes what the user values and plays a crucial role in the generation of Evoking strategies. Specific user goals are collected during the initialization (Step 0 in Figure 5). This is collected through the initial setup, see Sec. 4.2. The elements mentioned above are represented as word slots (enclosed in brackets in the input prompt), where the users’ actual context information can be inserted.
@@
(3) Prompt Optimization: To improve the GPT’s content quality and effectiveness, we carefully crafted the prompt according to OpenAI’s official guidelines (OpenAI, 2023). Sometimes LLMs can generate harmful, offensive, or biased texts (Zhuo et al., 2023; Hartvigsen et al., 2022). We employed an iterative prompt design process to ensure that the persuasive content is appropriate. Initially, one researcher created initial prompts and generated content using GPT-3.5, for six mental states (Figure 3) with five iterations each. Subsequently, two other researchers rated the satisfaction level of the generated content (scoring from 1 to 5), iterating until average satisfaction exceeded 4 to create content that is concise, appropriate, and engaging, while also aligned with our persuasive strategy. Through this process, we addressed some issues with LLM-generated content, such as gender-biased expressions and deviations from human preferences (like being overly exaggerated or stressful) by adding an additional ¡Notes¿ section to instruct LLMs’ generation. We also took steps to prevent hallucinations by avoiding certain real-world fact-related prompt statements that LLMs can easily make mistakes. By specifying the need to consider current activities in the ¡output format¿, we’ve made the outputs more contextually relevant. We note that this process cannot fully address the ethical concerns, which we further discuss in Sec. 8.3.3. Our final optimization prompt, shown in the middle top box in Figure 4, includes two modules: (a) The ¡Notes¿ module clarifies restrictions; (b) The ¡Output Format¿ module guarantees the correct output format.
@@
(4) Description of Persuasion Strategy: To ensure each generated content aligned with our proposed strategies, we provide a short description of each strategy based on our design in Sec. 3 and Figure 3, as indicated by the middle bottom box in Figure 4. Based on users’ in-the-moment mental states, we will select the corresponding strategies. Note that for the scaffolding habits strategy, we also input a habit selected from users’ initialization.
@@ 4.2. How to Intervene: Interaction Flow
Figure 5. Interaction Flow. Users first complete the global settings for their value and app list (Step 0). When opening a black-listed app, users need to first self-report their phone usage intent (Step 1). If the intent is habitual use, the app asks them to report their current mental state (Step 2). After that, a corresponding persuasion shows (Step 3).
Three figures about the intervention interface design. (Left) Two initialization pages. The left is about value setup, where there is a list of four value items and each item has two text edit boxes: goals, and habits. The value items are Career, Health, Life, and Hobbies. The right is about the blacklist of apps setup, where there is a list of apps installed on users’ phones, each accompanied by a checkbox. (Middle Left) The intent report intervention pop-up interface, and it has four buttons to choose from: habitual use, instrumental use, relax, I’m great and I will quit playing. (Middle Right) The mental state report intervention pop-up interface, and it has two single-choice questions: (1) Do you have any activity to do(”Yes” or ”No”), and (2) Do you have negative feelings at the moment, including options of ”Stress”, ”Boredom”, ”None”, or ”Other Negative Feelings”. The ”Other Negative Feelings” option has an edit text box next to it for description. (Right) Two persuasion pop-up interfaces. The top half is an interface for scaffolding habits and persuasion strategy. The interface is divided into three main sections. At the top, a row presents recommended habits, with an adjacent ’modify’ button for customization. In the center, a rectangular text box displays auto-generated persuasive content. The bottom right corner of this text box features ’thumb-up’ and ’thumb-down’ buttons for user feedback. At the bottom, there are two buttons labeled ’continue using’ and ’quit using,’ offering options to proceed or discontinue. The bottom half is an interface for understanding, comforting, and evoking strategy. This interface is similar to the previous one but lacks the top row of recommended habits. It still features the central text box with autogenerated persuasive content and ’thumb-up’ and ’thumb-down’ buttons at the bottom right. The options ’continue using’ and ’quit using’ remain at the bottom.
Figure 5. Interaction Flow. Users first complete the global settings for their value and app list (Step 0). When opening a black-listed app, users need to first self-report their phone usage intent (Step 1). If the intent is habitual use, the app asks them to report their current mental state (Step 2). After that, a corresponding persuasion shows (Step 3).
This section outlines the design of the interaction flow in our application. Our interaction process needs to achieve three functions: (1) collect users’ phone usage intent to identify habitual use, (2) obtain necessary information for prompt construction, and (3) display the generated content.
Following Takeaway \⃝raisebox{-0.9pt}{1} in Sec. 3.2, interventions should be targeted at habitual usage. Since automatic detection methods are unreliable (discussed further in Sec. 8), we ask users to self-report, and the system only triggers intervention when the user reports habitual use. In our prompt design in Sec. 4.1.2, we need two categories of information from users: their goals and habits, and their mental state. Goals and habits tend to be stable. So we integrated them into the app’s settings page, and users could adjust them as needed. Mental states, however, are more dynamic. Therefore, we captured them through participants’ self-reporting. In summary, each intervention episode includes three steps: (Step 1) the user reports their usage intent, (Step 2) the mental state, and (Step 3) the corresponding generated content is displayed, as shown in Fig. 5. Our final design is as follows:
@@
Step 0: Initialization. When initiating the MindShift app, we ask users to complete two global settings. The first is to set their values in four categories: career, health, life, and hobbies, detailing their goals and habits in each. This process serves two purposes: first, it provides the necessary goals for our Evoking strategy; second, it enables the creation of personalized habits in the Scaffolding Habits strategy. The second is to set a blacklist of apps. Launching apps from this list will trigger intervention.
@@
Step 1: Intent report. Interventions are only necessary when users habitually use their phones (Takeaway \⃝raisebox{-0.9pt}{1}). To collect users’ intents, we employ a self-reporting approach. Every time a blacklisted app is first opened during an unlock session, users choose from three options: ”Habitual use”, ”Instrumental use”, and ”Relaxation”. Only when users select ”Habitual use”, the intervention will proceed. Automatically detecting all use is beyond the scope of this paper. We envision that future work can automate this process, as discussed in Sec. 8.
@@
Step 2: Current mental state report. The mental state is a key factor that triggers users’ habitual phone use. Unlike physical context, detecting mental state is challenging due to the lack of mature techniques. Therefore, we ask users to self-report. Based on the mental states listed in Sec. 3.2 and Takeaway \⃝raisebox{-0.9pt}{3} & \⃝raisebox{-0.9pt}{4}, we propose two single-choice questions for users to report their mental state: (1) whether they are engaged in activities (”Yes” or ”No”), and (2) whether they currently have any negative feelings, including options of ”Stress”, ”Boredom”, ”None” (i.e., inertia), or ”Other Negative Feelings”. The ”Other Negative Feelings” option, with an adjacent text box for specifics, covers unlisted emotions. We tested these questions for reliability and validity (Price et al., 2015). To ensure validity (i.e., accurate reflection of mental states), three psychology experts verified the alignment of our questions with mental state coding in Sec. 3. To ensure reliability (answer consistency), we asked three pilot study participants to respond to situational mental state descriptions with two single-choice questions, achieving uniform responses. [^4] We plan to conduct more comprehensive validity and reliability testing in future work, as discussed in Sec. 8.
@@
Step 3: Persuasion. Based on our intervention content design in Figure 4, we leverage the power of GPT-3.5 to generate persuasive messages. The messages are displayed in a pop-up window, where users can choose to either quit the app or continue using it (the bottom of Step 3 in Figure 5). Additionally, users can also provide optional feedback by giving a thumbs up or down.
@@
Furthermore, for the Scaffolding Habits strategy, it is essential to link users’ own habits to specific use contexts. Therefore, we implement a user participation mechanism, adding an additional habit item along with an edit button (the upper interface of Step 3 in Figure 5). The habit item represents a system-generated suggestion based on the user’s current context and initial settings in Step 0. Users can edit and update their desired habits. Once submitted, the modified habit will be recommended the next time when users are in the same context.